How would you find the logarithmic or exponential equation of the graph, when one of the points is (4,1) and the other is at (1,0)?
Thanks!

Answer by ankor@dixie-net.com(12692)

How would you find the logarithmic or exponential equation of the graph, when one of the points is (4,1) and the other is at (1,0)?
:
This looks like a logarithmic function (x increases much faster than y)
:
It's easier to find the exponential function which is the inverse of this
Reverse the x,y values: 1,4 and 0,1
:
y = b^x
4 = b^1
b = 4
:
Check it using coordinates 0,1
y = 4^0
y = 1
:
the inverse of of an exponential function:
y = logb(x) and is the logarithmic function
:
y = log(base 4)(x)
x = 4^y
4 = 4^1
and
1 = 4^0